OpenStudy (anonymous):

What is the simplified form of each expression? 1/3(21m + 27) A. 63m + 9 B. 7m + 9 C. 7m + 81 D. 7m + 27

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } \times 21m + \frac{ 1 }{ 3 } \times 27\] Can you do it now?